OpenStudy (anonymous):

Tan15, I have to find the exact value using sum and difference formulas

OpenStudy (anonymous):

@Michele_Laino sorry to bother you but can you help with me with this problem?

OpenStudy (michele_laino):

we can write this: \[\Large \begin{gathered} \tan 15 = \tan \left( {45 - 30} \right) = \hfill \\ = \frac{{\tan 45 - \tan 30}}{{1 + \tan 45 \times \tan 30}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Yes then you get 1-(1/sqrt3)/1+(1)(1/sqrt3)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

Can you show me how to do the math for the next part

OpenStudy (michele_laino):

here are my steps: \[\Large \begin{gathered} \tan 15 = \tan \left( {45 - 30} \right) = \hfill \\ = \frac{{\tan 45 - \tan 30}}{{1 + \tan 45 \times \tan 30}} = \frac{{1 - \frac{1}{{\sqrt 3 }}}}{{1 + \frac{1}{{\sqrt 3 }}}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now the least common multiple at the denominator is sqrt(3), and the least common multiple at the denominator is also sqrt(39, so we can write: \[\Large \frac{{1 - \frac{1}{{\sqrt 3 }}}}{{1 + \frac{1}{{\sqrt 3 }}}} = \frac{{\frac{{\sqrt 3 - 1}}{{\sqrt 3 }}}}{{\frac{{\sqrt 3 + 1}}{{\sqrt 3 }}}} = \frac{{\sqrt 3 - 1}}{{\sqrt 3 }} \times \frac{{\sqrt 3 }}{{\sqrt 3 + 1}} = ...?\]

OpenStudy (anonymous):

Why did the 3rd piece come out as just sqrt 3-1? @Michele_Laino

OpenStudy (anonymous):

What step

OpenStudy (anonymous):

the sqrt 3-1/sqrt3

OpenStudy (michele_laino):

since I computed this: \[1 - \frac{1}{{\sqrt 3 }} = \frac{1}{1} - \frac{1}{{\sqrt 3 }} = \frac{{1 \times \sqrt 3 - 1}}{{\sqrt 3 }} = \frac{{\sqrt 3 - 1}}{{\sqrt 3 }}\]

OpenStudy (michele_laino):

namely I have computed the least common multiple between 1 and sqrt(3)

OpenStudy (michele_laino):

similarly it goes for denominator: \[1 + \frac{1}{{\sqrt 3 }} = \frac{1}{1} + \frac{1}{{\sqrt 3 }} = \frac{{1 \times \sqrt 3 + 1}}{{\sqrt 3 }} = \frac{{\sqrt 3 + 1}}{{\sqrt 3 }}\]

OpenStudy (michele_laino):

the last step is since we have to multiply the first fraction by the inverse of the second fraction